# EF-3DGS: Event-Aided Free-Trajectory 3D Gaussian Splatting

### v1

@@ Abstract
Scene reconstruction from casually captured videos has wide applications in real-world scenarios. With recent advancements in differentiable rendering techniques, several methods have attempted to simultaneously optimize scene representations (NeRF or 3DGS) and camera poses. Despite recent progress, existing methods relying on traditional camera input tend to fail in high-speed (or equivalently low-frame-rate) scenarios due to insufficient observations and large pixel displacements between adjacent frames. Event cameras, inspired by biological vision, record pixel-wise intensity changes asynchronously with high temporal resolution and low latency, providing valuable scene and motion information in blind inter-frame intervals. In this paper, we introduce the event camera to aid scene construction from a casually captured video for the first time, and propose Event-Aided Free-Trajectory 3DGS, called EF-3DGS, which seamlessly integrates the advantages of event cameras into 3DGS through three key components. First, we leverage the Event Generation Model (EGM) to fuse events and frames, supervising the rendered views observed by the event stream. Second, we adopt the Contrast Maximization (CMax) framework in a piece-wise manner to extract motion information by maximizing the contrast of the Image of Warped Events (IWE), thereby calibrating the estimated poses. Besides, based on the Linear Event Generation Model (LEGM), the brightness information encoded in the IWE is also utilized to constrain the 3DGS in the gradient domain. Third, to mitigate the absence of color information of events, we introduce photometric bundle adjustment (PBA) to ensure view consistency across events and frames. Additionally, we propose a Fixed-GS training strategy that separates the optimization of scene structure and color, effectively addressing color distortions caused by the lack of color information in events. We evaluate our method on the public Tanks and Temples benchmark and a newly collected real-world dataset, RealEv-DAVIS. Our method achieves up to 2dB higher PSNR and 40% lower Absolute Trajectory Error (ATE) compared to state-of-the-art methods under challenging high-speed scenarios. Our project page is https://lbh666.github.io/ef-3dgs/.
Index Terms:
Event Camera, Novel View Synthesis, 3D Gaussian Splatting, Neural Rendering
@@
Figure 1: Free-trajectory 3DGS under high speed. The top row shows the overall paradigm. The colored dots (red for positive and blue for negative events) in the top row represent the event data. Given free-trajectory video and corresponding event stream under the high-speed scenario, we aim to jointly estimate poses and the 3DGS. Our method surpasses current state-of-the-art methods in terms of both rendered results (middle row) and pose estimation (bottom row).
I Introduction
@@
In recent years, Neural Radiance Fields (NeRF) [1, 2, 3] and 3D Gaussian splatting (3DGS) [4, 5] have made significant progress in novel view synthesis tasks. Given a set of posed images of the same scene, they optimize an implicit or explicit scene representation using volume rendering. While subsequent methods [2, 3, 6, 7] have made great strides in rendering quality and speed, how to reconstruct the scene from a video captured with free camera trajectories remains an open problem. It has great application potential in the real world, such as in VR/AR, video stabilization, and mapping. To tackle this challenging task, several efforts have been made.
Free-trajectory videos often cover a vast field of view, so a key challenge is how to preserve the detailed textures of the foreground while representing the overall large-scale scene. Previous methods [8, 9] adopt different spatial transformation and allocation approaches based on neighboring frames’ viewing angles and range to achieve efficient and high-quality scene reconstruction. Another challenge is that accurate pose estimation is often difficult to obtain in free-trajectory scenarios. Previous methods [9, 10] have incorporated priors, such as the depth [11, 12, 13] and the optical flow [14], into their loss functions to improve pose estimation or modify their optimization strategies. While these methods above can render photo-realistic images in typical free-trajectory scenarios, the rendering quality degrades significantly in high-speed scenarios (or equivalent low-frame-rate scenarios). This type of scenario has essential applications in the real world, such as autonomous driving or FPV exploration. Besides, the significant motion between consecutive frames poses challenges for feature matching and further abrupt pose estimation (as shown in the third row of Fig.1). This makes it even more crucial for the algorithm to robustly estimate the poses in such scenarios.
@@
The performance degradation of prior methods can be attributed to two primary factors. Firstly, the limited number of camera observations leads to an under-constrained scene reconstruction problem. This can cause the scene representation to converge to a trivial solution [15, 16, 17], where the model overfits to the training views without capturing the correct underlying geometry structure. Consequently, the synthesized images from novel viewpoints exhibit visual artifacts. Secondly, the substantial discrepancies between consecutive frames, resulting in diminished overlapping regions, violate the implicit assumption of continuous motion between adjacent frames, which was leveraged by previous techniques such as LocalRF [9] and NoPe-NeRF [10]. This significant violation greatly exacerbates the ill-posedness of the joint optimization of scene and camera poses, further deteriorating the quality of the synthesized novel views.
@@
Event camera is a bio-inspired image sensor that asynchronously records per-pixel brightness changes, offering advantages such as high temporal resolution, high dynamic range, and no motion blur [18, 19, 20, 21, 22, 23]. The brightness information recorded in the event stream can effectively complement the missing scene information between consecutive frames. Moreover, in scenarios with stable scene illumination, the event data naturally encodes the motion information of the scene [24, 25, 26], containing rich motion cues. These properties make event cameras well-suited for scene reconstruction tasks in high-speed and free-trajectory scenarios [27]. However, seamlessly integrating the aforementioned benefits of event cameras is nontrivial. First, NeRF or 3DGS renders absolute pixel brightness, which aligns with image data. In contrast, event stream records sparse scene brightness changes, lacking the ability to perceive absolute brightness, and most event cameras lack color information. Second, the motion cues in event cameras is encoded in the spatio-temporal correlations of the event stream. In contrast, previous methods typically sample and optimize frames and pixels independently, without explicitly considering these temporal correlations.
@@
In this work, we propose Event-Aided Free-Trajectory 3DGS, dubbed EF-3DGS, a framework that integrates event data into the scene optimization process to fully leverage the high temporal resolution property of event cameras. Specifically, based on the Event Generation Model (GEM), we introduce an event-based re-render loss, which extends the 3DGS optimization to the continuous event stream. This allows us to effectively utilize the brightness cues encoded in the event streams between adjacent frames, providing rich supervisory signals to alleviate the problem of insufficient sparse viewpoints. Regarding the pose estimation challenge, we introduce the CMax [28] framework to exploit the spatio-temporal correlations of events. We obtain the motion field by leveraging the pseudo-depth from 3DGS rendering and the relative camera motion between consecutive time. We then warp the events triggered by the same edge along the motion trajectories to maximize the sharpness of the image of warped events (IWE), thereby estimating the motion that best matches the current spatio-temporal event patterns. Furthermore, through the Linear Event Generation Model (LEGM) [29, 30], we establish a connection between motion and brightness changes. This allows us to constrain the 3DGS in the gradient domain using the IWE. Finally, as event data primarily records scene brightness changes, lacking color information, we introduce photometric bundle adjustment (PBA) and a Fixed-GS strategy to address this. PBA recovers color by optimizing reprojection errors onto RGB frames, while Fixed-GS enables separate optimization of scene structure and color.
@@
  • Experiments on both public benchmarks and real-world datasets demonstrate that our method significantly outperforms existing state-of-the-art approaches in terms of both rendering quality and trajectory estimation accuracy. Our method achieves up to 2dB higher PSNR and 40% lower Absolute Trajectory Error compared to state-of-the-art methods in high-speed scenarios.
II RELATED WORKS
II-A Large Scale Scene Reconstruction
The reconstruction of large scenes has emerged as a significant area of research over the past decade. Traditional approaches [31, 32] rely on Structure from Motion techniques to recover camera poses and sparse 3D point clouds. The introduction of NeRF [1] marked a paradigm shift in scene representation and reconstruction. Several works [33, 34, 2] have attempted to extend NeRF to unbounded scenes. F2-NeRF [8] adopts perspective warping based on the local viewing range to handle free-viewpoint unbounded scenes. To extend unbounded scenes to large-scale scenes, Mega-NeRF [35] and Block-NeRF [36] adopt a divide-and-conquer strategy. LocalRF [9] allocates new radiance fields dynamically according to the distance between the current view and the previous radiance field. Recently, 3D Gaussian Splatting (3DGS) has been extended to large-scale scene reconstruction due to its efficient rendering capabilities and explicit 3D representation. CityGaussian [37] and VastGaussian [38] adopt spatial allocation strategies similar to Mega-NeRF [35]. CityGaussian [37] dynamically selects poses for optimization based on visibility criteria, while VastGaussian [38] employs a visibility-driven progressive approach to dynamically select both training poses and Gaussian points for optimization. In contrast to these methods, which primarily focus on scalability and often require precomputed camera poses from COLMAP [39], we focus on the challenges of reconstructing scene from casually captured free-trajectory videos without relying on precomputed poses.
II-B Joint Pose and Scene Optimization
@@
The research community has recently focused on developing methods [17, 9, 40, 41, 10, 42, 5] that can be optimized without requiring precomputed camera poses. A line of work has focused on improving the stability of the optimization process. GARF [41] and BARF [42] both find that the high-frequency position encoding is prone to local minima and try to improve it. For example, GARF [41] proposes using Gaussian activation to replace the sinusoidal position encoding. Another line of work has investigated incorporating additional constraints to make the problem more tractable. SPARF [17] incorporates cross-view feature matching constraints and depth consistency regularization. More recent approaches [10, 9, 5] leverage pre-trained networks, i.e., monocular depth estimation and optical flow estimation, to obtain additional priors and better address the pose-free NeRF reconstruction problem. Exploiting 3DGS’s explicit representation, CF-3DGS [5] directly back-projects Gaussian points using depth maps, enhancing the utilization of depth priors. While the aforementioned methods have made notable progress, they have yet to fully address the challenges posed by high-speed ( or equivalently sparse-view) scenarios or rely on a good pose initialization. Our approach addresses these issues by leveraging motion and brightness cues from event streams.
II-C Event-Based Novel view Synthesis
@@
Recent works have explored the integration of event cameras into the NeRF framework. Early approaches, such as E-NeRF [43] and EventNeRF [44], utilize event-based generative models, minimizing the difference between the rendered brightness changes and observed brightness changes calculated by the accumulated events. Building upon this, Robust e-NeRF [45] incorporates a more realistic imaging model into the event-based framework, accounting for factors like refractory periods and noise. Beyond event-based NeRF, efforts have also been made to integrate event data into image-based methods. For instance, E2NeRF [46] and EvDeblurNeRF [47] leverage the Event Double Integral (EDI) [48] model to address the deblurring problem, while DE-NeRF [49] and EvDNeRF [50] leverage the high temporal resolution property of event camera to capture fast-moving elements in dynamic scene. A key distinction of our work is that, unlike the prior methods that rely on accurate precomputed poses, we jointly optimize for both the camera poses and the scene representation. Furthermore, while previous works have been limited to simulated environments or simple indoor scenes, we evaluate our approach in large-scale outdoor scenarios.
III Preliminary
@@
3DGS [4] parametrizes the 3D scene as a set of 3D gaussians $\{G_{k}\}_{k=1}^{K}$ that carry the geometric and appearance information. Each 3D Gaussian is characterized by several learnable properties, including its center position $\mu\in\mathbb{R}^{3}$, opacity $\alpha\in[0,1]$, spherical harmonics (SH) features $\mathbf{f}_{k}\in\mathbb{R}^{3\times 16}$ for view-dependent color $c\in\mathbb{R}^{3}$, rotation matrix $R\in\mathbb{R}^{3\times 3}$ (stored in quaternion form), scale factor $s\in\mathbb{R}^{3}$. The shape of each Gaussian is defined by the covariance matrix $\mathbf{\Sigma}$ and the center (mean) point $\mu$,
$G(x)=\exp(-\frac{1}{2}{(x-\mu)}^{T}\mathbf{\Sigma}^{-1}(x-\mu)).$ (1)
To preserve the valid positive semi-definite property during optimization, the $\mathbf{\Sigma}$ is derived by the scale factor $s$ and the rotation factor $R$,
$\mathbf{\Sigma}=RSS^{T}R^{T}.$ (2)
Before rendering, the Gaussian is first projected to the 2D image plane:
$\mathbf{\Sigma}^{2D}=JW\mathbf{\Sigma}W^{T}J^{T},$ (3)
where J is the Jacobian of the affine approximation of the projective transformation, and W is the viewing transformation. During rendering, a tile-based rasterizer is applied to enable fast sorting and $\alpha$-blending. The color of each pixel is calculated via blending N ordered overlapping points:
@@
where $c_{i}$ is calculated from spherical harmonics and view direction, $\alpha_{i}$ is the multiplication of opacity and the transformed 2D Gaussian and r denotes the image pixel. With the forward rendering procedure, we can optimize 3DGS by minimizing the photometric loss between observation and rendered pixels:
$\mathcal{L}_{color}=(1-\lambda)\mathcal{L}_{1}(\hat{I},I)+\lambda\mathcal{L}_{D-SSIM}(\hat{I},I),$ (5)
where the first term is a standard $\mathcal{L}_{1}$ loss between the target and rendered images, $\mathcal{L}_{D-SSIM}$ is a differentiable version of structural similarity index and $\lambda$ is balancing weight which is set to 0.2 following [4]. As this optimization is highly sensitive to local minima, 3DGS additionally employs periodic adaptive densification and pruning of splats. By integrating depth $t_{i}$ of each Gaussian in Equation 4 along the ray, we can also obtain a expected depth value $\hat{D}(\mathbf{r})$:
$\hat{D}(\mathbf{r})=\sum_{i=1}^{N}t_{i}\alpha_{i}\ \prod_{j=1}^{i-1}(1-\alpha_{j}).$ (6)
Figure 2: Method Overview. Our method takes video frames $\{I_{i}\}$ and event stream $\varepsilon=\{\mathbf{e}_{k}\}$ as input. During training, we random sample $t\in\{t_{i,j}\}$, leveraging the events within current subinterval and most recent frame to establish the proposed three main constrains,$\mathcal{L}_{EGM}$, $\mathcal{L}_{LEGM}$ and $\mathcal{L}_{PBA}$. The colored dots (red for positive and blue for negative events) represent the event data. The details of $\mathcal{L}_{LEGM}$ and CMax framework are illustrated in Fig. 3
IV Method
In this section, we introduce our method. The overall framework is shown in Fig. 2. Given a video of a free-trajectory $\{I_{i}\}$ captured at time $\{t_{i}\}$ and an event stream $\varepsilon=\{\mathbf{e}_{k}\}$, our goal is to reconstruct the 3DGS of the scene and the corresponding camera trajectory$\{T_{i}\}$. The photometric loss of the original 3DGS follows an analysis-by-synthesis process. Similarly, in order to incorporate the events into the 3DGS optimization process, we delve into the imaging principles of event cameras. Revolving around the Event Generative Model (EGM) and Linearized Event Generative Model (LEGM), we establish constraints from the connection between motion and brightness changes. Further, through photometric bundle adjustment (PBA) and Fix-GS training strategy, we mitigate the absence of color information of events and further strengthen the consistency of cross-view rendering.
IV-A EGM Driven Optimization
The Event Generation Model (EGM) describes the base working principle of event cameras, which capture pixel-wise brightness changes asynchronously. When the logarithmic brightness change at a pixel $\mathbf{u}_{k}=(x_{k},y_{k})$, exceeds a predefined contrast threshold $C$,

### v3

@@ Abstract
Scene reconstruction from casually captured videos has wide real-world applications. Despite recent progress, existing methods relying on traditional cameras tend to fail in high-speed scenarios due to insufficient observations and inaccurate pose estimation. Event cameras, inspired by biological vision, record pixel-wise intensity changes asynchronously with high temporal resolution and low latency, providing valuable scene and motion information in blind inter-frame intervals. In this paper, we introduce the event cameras to aid scene construction from a casually captured video for the first time, and propose Event-Aided Free-Trajectory 3DGS, called EF-3DGS, which seamlessly integrates the advantages of event cameras into 3DGS through three key components. First, we leverage the Event Generation Model (EGM) to fuse events and frames, enabling continuous supervision between discrete frames. Second, we extract motion information through Contrast Maximization (CMax) of warped events, which calibrates camera poses and provides gradient-domain constraints for 3DGS. Third, to address the absence of color information in events, we combine photometric bundle adjustment (PBA) with a Fixed-GS training strategy that separates structure and color optimization, effectively ensuring color consistency across different views. We evaluate our method on the public Tanks and Temples benchmark and a newly collected real-world dataset, RealEv-DAVIS. Our method achieves up to 3dB higher PSNR and 40% lower Absolute Trajectory Error (ATE) compared to state-of-the-art methods under challenging high-speed scenarios.
@@
Figure 1: Free-trajectory 3DGS under high speed. The top row shows the overall paradigm. The colored dots (red for positive and blue for negative events) in the top row represent the event data. We leverage continuous event streams to aid discrete video frames captured along free trajectories in high-speed scenarios, jointly optimizing camera poses and reconstructing the 3DGS. Our method surpasses current state-of-the-art methods in terms of both rendered results (middle row) and pose estimation (bottom row).
1 Introduction
@@
In recent years, Neural Radiance Fields (NeRF) [30, 1, 31] and 3D Gaussian splatting (3DGS) [20, 9] have made significant progress in novel view synthesis tasks. Given a set of posed images of the same scene, they optimize an implicit or explicit scene representation using volume rendering. While subsequent methods [1, 31, 7, 2] excel with posed images, reconstructing scenes from videos with free camera trajectories remains challenging despite its applications in VR/AR, video stabilization, and mapping. To tackle this challenging task, several efforts have been made.
Accurate pose estimation is often difficult to obtain in free-trajectory scenarios, which directly impacts the quality of scene reconstruction. One line of work draws inspiration from Simultaneous Localization and Mapping (SLAM). They [28, 9, 19] follow its optimization paradigm, progressively optimizing camera trajectories and alternating between camera pose and scene refinement. Another line of work [4, 19, 25, 9, 42] explores incorporating additional geometric or motion priors such as depth estimation [29, 36] or optical flow [41] to establish constraints beyond photometric rendering loss. While these methods can render photo-realistic images in typical free-trajectory scenarios, both their rendering quality and pose estimation accuracy degrade significantly in high-speed scenarios (or equivalently low-frame-rate scenarios) as shown in Fig. 1. Such high-speed scenarios have essential applications such as autonomous driving and First-Person View (FPV) exploration.
@@
The performance degradation of prior methods can be attributed to two primary factors. First, the limited number of camera observations leads to an under-constrained scene reconstruction problem. This can cause the scene representation to converge to a trivial solution [32, 47, 42], where the model overfits to the training views without capturing the correct underlying geometry structure. Second, the substantial discrepancies between consecutive frames, resulting in diminished overlapping regions, violate the implicit assumption of continuous motion between adjacent frames, which is leveraged by previous methods. Moreover, geometric and motion priors like optical flow and feature matching become unreliable in such scenarios. These significant violations greatly exacerbate the ill-posedness of the joint optimization of scene and camera poses.
@@
Event camera is a bio-inspired image sensor that asynchronously records per-pixel brightness changes, offering advantages such as high temporal resolution, high dynamic range, and no motion blur [12, 17, 45, 46, 52, 6]. The brightness information recorded in the event stream can effectively complement the missing scene information between consecutive frames. Moreover, the event data naturally encodes the motion information of the scene [24, 40, 44], containing rich motion cues. These properties make event cameras well-suited for scene reconstruction tasks in high-speed and free-trajectory scenarios. However, seamlessly integrating the aforementioned benefits of event cameras is nontrivial. First, 3DGS renders absolute pixel brightness, which aligns with image data. Event cameras, however, record sparse differential brightness changes. Directly integrating the differential operations into 3DGS may amplify noise and lead to ill-conditioned optimization problems with high sensitivity to parameter initialization and perturbations. Second, event cameras encode motion through continuous spatio-temporal trajectories of events. In contrast, frame-based data inherently discretizes continuous motion, forcing traditional methods to rely on correspondence matching, which fails in high-speed scenarios with large inter-frame displacements. These fundamental challenges require carefully designed method that bridges the gap between the event data and 3DGS optimization.
@@
In this work, we propose Event-Aided Free-Trajectory 3DGS, dubbed EF-3DGS, a framework that integrates event data into the scene optimization process to fully leverage its high temporal resolution property. Our approach comprises three key components: (1) In the Event Generation Model (EGM), we introduce an event-based re-render loss, which extends the 3DGS optimization to the continuous event stream. This allows us to utilize the brightness cues encoded in the event stream between adjacent frames, providing rich supervisory signals to alleviate the insufficient sparse view issues. (2) In the Linear Event Generation Model (LEGM), regarding the pose estimation challenge, we introduce the CMax [11] framework to exploit the spatio-temporal correlations of events. We obtain the motion field by leveraging the pseudo-depth from 3DGS rendering and the relative camera motion between consecutive frames. We then warp the events triggered by the same edge along the motion trajectories to maximize the sharpness of the image of warped events (IWE), thereby estimating the motion that best matches the current spatio-temporal event patterns. Furthermore, through the Linear Event Generation Model [10, 13], we establish a connection between motion and brightness changes. This allows us to constrain the 3DGS in the gradient domain using the IWE. (3) As most event data primarily records scene brightness changes, lacking color information, we introduce photometric bundle adjustment (PBA) and a Fixed-GS strategy to address this. PBA recovers color by optimizing reprojection errors onto RGB frames, while Fixed-GS enables separate optimization of scene structure and color.
@@
  • Experiments on both public benchmarks and real-world datasets demonstrate that our method significantly outperforms existing state-of-the-art approaches in terms of both rendering quality and trajectory estimation accuracy.
2 Related Works
2.1 Joint Pose and Scene Optimization
@@
The research community has recently focused on developing methods [42, 28, 50, 8, 4, 23, 9] that can be optimized without requiring precomputed camera poses. A line of work has focused on improving the stability of the optimization process. GARF [8] and BARF [23] both find that the high-frequency position encoding is prone to local minima and try to improve it. For example, GARF [8] proposes using Gaussian activation to replace the sinusoidal position encoding. Another line of work has investigated incorporating additional constraints to make the problem more tractable. LocalRF [28] leverages the prior assumption of continuous motion between adjacent frames and progressively adds and optimizes camera poses. More recent approaches [4, 28, 9] leverage pre-trained networks, i.e., monocular depth estimation and optical flow estimation. Exploiting 3DGS’s explicit representation, CF-3DGS [9] directly back-projects Gaussian points using depth maps. While the aforementioned methods have made notable progress, they have yet to fully address the challenges posed by high-speed scenarios or rely on a good pose initialization. Our approach addresses these issues by leveraging motion and brightness cues from event streams.
2.2 Event-Based Novel View Synthesis
@@
Recent works have explored the integration of event cameras into the NeRF or 3DGS framework. Early approaches, such as E-NeRF [21] and EventNeRF [37], utilize event-based generative models, minimizing the difference between the rendered brightness changes and observed brightness changes. Building upon this, Robust e-NeRF [26] incorporates a more realistic imaging model into the event-based framework, accounting for factors like refractory periods and noise. Beyond event-based NeRF, efforts have also been made to integrate event data into image-based methods. For instance, E2NeRF [35] and EvDeblurNeRF [5] leverage the Event Double Integral (EDI) [33] model to address the deblurring problem, while DE-NeRF [27] and EvDNeRF [3] leverage the high temporal resolution property of event cameras to capture fast-moving elements in dynamic scene. More recently, Event-3DGS [16] and EaDeblur-GS [51] have extended previous approaches to 3D Gaussian Splatting, achieving superior rendering quality and real-time performance. A key distinction of our work is that, unlike the prior methods that rely on accurate precomputed poses, we target free-trajectory scenarios, jointly optimizing for both the camera poses and the scene representation. Furthermore, while previous works have been limited to simulated and simple environments, we evaluate our approach in large-scale outdoor scenarios with complex motions and lighting conditions.
Figure 2: Method overview. Our method takes video frames and event stream as input. During the first stage, we progressively add new event images, leveraging the events and most recent frame to establish the event-driven optimization. During the second stage, we adopt the Fixed-GS strategy to mitigate the color distortion of 3DGS. The details of $\mathcal{L}_{LEGM}$ and CMax framework are illustrated in Fig. 3.
3 Preliminary
@@
3DGS [20] parametrizes the 3D scene as a set of 3D gaussians $\{G_{k}\}_{k=1}^{K}$ that carry the geometric and appearance information. Each 3D Gaussian is characterized by several learnable properties, including its center position $\mu\in\mathbb{R}^{3}$, opacity $\alpha\in[0,1]$, spherical harmonics (SH) features $\mathbf{f}_{k}\in\mathbb{R}^{3\times 16}$ for view-dependent color $c\in\mathbb{R}^{3}$, rotation matrix $R\in\mathbb{R}^{3\times 3}$ (stored in quaternion form), scale factor $s\in\mathbb{R}^{3}$. The shape of each Gaussian is defined by the covariance matrix $\mathbf{\Sigma}$ and the center (mean) point $\mu$, $G(x)=\exp(-\frac{1}{2}{(x-\mu)}^{T}\mathbf{\Sigma}^{-1}(x-\mu)).$ During rendering, a tile-based rasterizer is applied to enable fast sorting and $\alpha$-blending. The color of each pixel is calculated via blending N ordered overlapping points:
@@
where $c_{i}$ is calculated from spherical harmonics and view direction, $\alpha_{i}$ is the multiplication of opacity and the transformed 2D Gaussian and r denotes the image pixel. With the forward rendering procedure, we can optimize 3DGS by minimizing a weighted combination loss of $\mathcal{L}_{1}$ and $\mathcal{L}_{D-SSIM}$ between observation and rendered pixels: $\mathcal{L}_{color}=(1-\lambda)\mathcal{L}_{1}(\hat{I},I)+\lambda\mathcal{L}_{D-SSIM}(\hat{I},I),$ where $\lambda$ is balancing weight which is set to 0.2 following [20]. By integrating depth $d_{i}$ in Equation 1 along the ray, we can also obtain a expected depth value $\hat{D}(\mathbf{r})$:
$\hat{D}(\mathbf{r})=\sum_{i=1}^{N}d_{i}\alpha_{i}\ \prod_{j=1}^{i-1}(1-\alpha_{j}).$ (2)
4 Method
The overall framework is shown in Fig. 2. Given a video of a free-trajectory $\{I_{i}\}$ captured at time $\{t_{i}\}$ and the event stream $\varepsilon=\{\mathbf{e}_{k}\}$, our goal is to reconstruct the 3DGS of the scene and the corresponding camera trajectory$\{T_{i}\}$. Following the analysis-by-synthesis paradigm of 3DGS, we extend this approach by incorporating event camera data through two fundamental imaging principles: the Event Generative Model (EGM) and Linear Event Generative Model (LEGM). To address the absence of color information in events and ensure cross-view consistency, we further introduce photometric bundle adjustment (PBA) and a Fixed-GS training strategy.
4.1 EGM Driven Optimization
The Event Generation Model (EGM) describes how event cameras asynchronously record pixel-wise brightness changes. When the logarithmic brightness change at a pixel $\mathbf{u}_{k}=(x_{k},y_{k})$, exceeds a predefined contrast threshold $C$,